If' ‘g' is the acceleration due to gravity on the earth's surface, the gain in the potential energy of an object of mass m raised from the surface of the earth. to a height equal to the radius ‘R’ of the earth is
 
(a) ½ mgR
(b) ½ mgR
(c) 2 mgR
(d) mgR

Grade:6

3 Answers

Deepak Patra
9 years ago

(b) At earth surface, P.E. of system is – GmM/R At a distance R from the earth’s surface, P.E. of system is – GmM/2R

? ? U = - GmM / 2R + GmM/R ; ?U = GmM /2R

Now GM/R2 = g; ? GM / R = gR ? ?U = ½ mgR
Raunaq Mehta
7 years ago
(b) At earth surface, P.E. of system is –GmM/R At a distance R from the earth’s surface, P.E. of system is – GmM/2R? ? U = - GmM / 2R + GmM/R ; ?U = GmM /2RNow GM/R2 = g; ?GM / R = gR ? ?U = ½ mgR
ankit singh
3 years ago
Gravitational potential energy on the earth surface, Ur = -GMmR Gravitational potential energy at a height h ... If g is the acceleration due to gravity on the earth's surface, the gain in the potential energy of an object of mass m raised from the surface of the earth to a height equal to the radius R of the earth is.